OpenStudy (anonymous):

A pharmacist mixed some 10%-saline solution with some 15%-saline solution to obtain 100 mL of a 12%-saline solution. How much of the 10%-saline solution did the pharmacist use in the mixture? A 60 mL B 45 mL C 40 mL D 25 mL

OpenStudy (anonymous):

\[ 10x + 15( 100 - x ) = 12 (100 )\\\\ -5 x = 1200- 1500 \\\\ x = 60 \ ml\]